OpenStudy (anonymous):

Use elimination to solve the system of equations. x + 3y = 13 -2x + 6y = 22 a.(4,2) b.(1,4) c(1,3) d.(4,3)

OpenStudy (anonymous):

You want to get rid of one of the variables, so multiply the top equation by two and add them together. 12y=48, solve for y and plug it into one of the others to find the answer

OpenStudy (anonymous):

2x + 6y = 26 -2x + 6y = 22 12y = 48 y = 4 solve for x 2x +6(4) = 26 2x = 26 - 24 2x = 2 x = 1 hope i answered ir qn

OpenStudy (anonymous):

b is the answer

OpenStudy (anonymous):

Thanks Friend :) Thanks for the answer and demostration

OpenStudy (anonymous):

pleasure